OpenStudy (anonymous):

2sqrt6 + 4 + sqrt6 sqrt=square root

OpenStudy (anonymous):

\[2\sqrt{6}+4+\sqrt{6}=3\sqrt{6}+4\]

OpenStudy (anonymous):

You just collect your sqrt(6)'s since you can't simplify them any further.

OpenStudy (anonymous):

Thanks Lokisan:)

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

You're a lifesaver:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

So what does Lokisan stand for?

OpenStudy (anonymous):

Nothing much. Just punched it in and it sounded good (to me, anyway)!